Using divisibility tests, determine which of the following numbers are divisible by 11: 
10824

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Sum of digits at odd places = 4 + 8 + 1 = 13
Sum of digits at even places = 2 + 0 =2
Difference = 13–2 = 11
Since, the difference between sum of digits at odd places and sum of digits at even places is 11 which is divisible by 11.
Hence, 10824 is divisible by 11

Thanks